# Jon Danielsson London School of Economics Andreas Uthemann

### 10 Jan 2024 v3

@@ 3.3 Six criteria for AI use in financial policy
4 AI use in the financial system
Many public and private financial system applications can be supported by the two forms of AI accessible today, reactive and limited memory. And, as the engines’ capabilities develop at a rapid pace, and we come closer to a theory of mind, and possibly AGI, we expect AI use to expand in both the public and private sectors (Maple et al., 2023; Kinywamaghana and Steffen, 2021).
Credit allocation, fraud, compliance and risk management are just a few of the tasks in which reactive machines have proven highly valuable. Supervisors may use them to develop the rulebook and enforce compliance with it, as the UK FCA is increasingly doing. There is no technical reason why AI cannot perform the majority of such jobs now; what is holding it back is a lack of human expertise, compute and data sharing (see Section 2.1) as well as caution.
For private financial institutions, where senior managers can be held to account, the unknown territory of AI, compared to relatively safe and familiar incumbent technology, might represent unacceptable personal risk. However, while traditional financial institutions might be held back by legacy systems, and perhaps a culture that reflects traditional practices, challenger banks can find it much easier to use technology. In a competitive market, their greater efficiency and better delivery of financial services will most likely push AI adoption across the private sector.
The authorities have chosen a more cautious approach to AI. However, as the private sector develops its use of AI, the authorities must follow suit if they are to remain relevant. They are already doing so, as evidenced by Moufakkir (2023) and Cook (2023). Anecdotal evidence suggestes that current AI use is mostly limited to regular forecasting and low-level data processing. However, most authorities are only now beginning a more structured analysis of AI.
4.1 Authorities taking advantage of AI
While financial authorities have so far used AI fairly sparingly, we expect that to change. This will necessitate investments in data, human capital and computing power. The most straightforward is data, as authorities already have access to large databases, including confidential disclosures, that can be used to train AI. This includes the following:
  • Observations on past compliance and supervisory decisions
  • Prices, trading volumes, and securities holdings fixed-income, repo, derivatives and equity markets
  • Assets and liabilities of commercial banks
  • Network connections, like cross institution exposures, including cross border
  • Textual data
    • The rulebook
    • Central bank speeches, policy decisions, staff analysis
    • Records of past crisis resolution
  • Internal economic models
    • Interest rate term structure models
    • Market liquidity models
    • Inflation, GDP and labour market forecasting models
    • Equilibrium macro model for policy analysis
The authorities will find human resources and compute more difficult, not the least because of costs, as we discuss below in Section 5.4. Perhaps the most obvious is the cost of human capital specialised in AI. Furthermore, training AI engines is very expensive and necessitates access to specialised facilities, both of which are out of reach for most financial authorities. However, because many of the largest costs are associated with training general-purpose engines, it is easy to overstate those problems. The marginal cost to the authorities is likely to be significantly smaller, as they will may need to extend extant engines via transfer learning on curated dataset for financial sector use that include the items listed above.
There are many activities within the financial authorities where AI could be of considerable help. Start with the microprudential regulations, both design and supervision. ML can be used to translate the rulebook into a logic engine, allowing analysis of internal consistency and coverage and facilitating AI supervision. Several authorities, like the UK’s FCA, have already made progress in this area. Generative models can simulate undesirable behaviour and hence help in designing the rulebook to protect against such outcomes. Reactive engines can identify undesirable behaviour not covered by extant rules and propose remedies. AI can furthermore be beneficial in supervision, providing routine compliance advice to the private sector, recommending regulatory actions and even making supervisory decisions. While human supervisors would initially closely oversee it, reinforcement learning with human feedback will likely help supervisory AI to become increasingly performant. For challenging areas of microprudential supervision, adversarial architecures such as generative adversarial networks (GANs) might be particularly beneficial in understanding complex areas of authority-private sector interactions, such as fraud detection.
It will also likely be useful in other activities such as ordinary economic analysis and forecasting thanks to an economic engine that performs both theoretical and statistical analysis. This could be achieved with a general purpose foundation model augmented via transfer learning using publicly and privately available data, established economic theory embodied in economic models and previous policy analysis. Reinforcement learning with feedback from human experts might be useful to further improve the engine. Such AI should furthermore be very beneficial to those conducting macroprudential stress tests. A key challenge in such analysis (Anderson et al., 2018) is that the stress needs to cover behaviour that we don’t usually see in the system and capture the particular interrelations between the various market participants in times of stress. Here generative models could be particularly valuable in facilitating the running of scenarios in crisis resolution, helping to identify and analyse the drivers of extreme market stress. The authorities could also use such models as artificial labs to experiment on policies and evaluate private sector algorithms, perhaps building on existing agent-based simulations. Over time, as AI technology improves, theory of mind AI might be able to understand and provide context for policy decisions and hence be invaluable in crisis resolution. AI that understands how its interlocutors think can explain its decisions to them, as well as reason about how they will react to its actions and plan accordingly, will be of considerable benefit. Not the least, it could address the problem of misspecified objectives.
4.2 The risk of authorities losing control
While an authority might not wish to get to that point, its use of AI might end up there regardless. As we come to trust AI analysis and decisions and see how well it performs in increasingly complex and essential tasks, it may end up by stealth, where the authorities do not want to be. By then, turning the AI engine off may be impossible because it performs vital functions, and the risk of disastrous outcomes might be deemed unacceptably high. Eventually, we risk becoming dependent on a system for critical analysis and decisions we don’t entirely, or even partially, understand.
@@ 5 The channels for how AI can destabilise the financial system
While we anticipate that AI will be broadly beneficial to the financial system, it may also increase systemic financial risk, defined as a serious financial crisis that has a high chance of causing an economic recession. The costs of systemic crises are very high, in the trillions of dollars for large economies, as noted by Barnichon et al. (2022). As such crises are usually caused by excessive low-quality lending, amplified by high leverage and asymmetric information, it should be simple to prevent them, which is the object of macroprudential regulations. It is not so in reality. The almost infinite complexity of the financial system makes it very hard to design effective regulations that simultaneously prevent excesses and do not hamper economic growth too much. On top of that, because the economy is usually growing rapidly before a crisis, there is strong political opposition to any measures that might threaten that growth.
@@
We build on extant work on AI safety, like Weidinger et al. (2022), Bengio et al. (2023) and Shevlane et al. (2023), which identifies several societal risks arising from AI use, including malicious use, misinformation and loss of human control. When augmenting that, we propose five channels for how AI use may destabilise the financial system, all of which follow from the AI issues discussed in Section 3 and the conceptual challenges in Section 2.
@@ 5.1 Malicious use of AI
The first channel arises from how the operators of financial AI can use it to optimise against the system for malicious purposes. Financial institutions and employees working for them deploy significant financial resources in their pursuit of profit and usually are not very concerned about the wider social impact of their activities. They can change the system in ways that benefit them while not being detectable by others, as we have seen many times in the past. AI provides such actors with new opportunities, either via adversarial attacks by feeding particular data into its training algorithms or influencing network design. That might not even be necessary since it would be straightforward to take advantage of the inability to specify all the objectives AI has to meet to exploit gaps in how AI sees its responsibilities.
One way such malicious use might occur is by manipulating the governance process for conduct in the financial system. The financial authorities and internal control aim to align the interests of system participants to the organisation that employs them and society. AI is particularly effective in finding profitable loopholes and amplifying vulnerabilities, facilitating misbehaviour that, while legal, is damaging both to society and even the institution employing the AI.
The operator of the AI could take that one step further and use it to deliberately create market stress, which, of course, is profitable to those forewarned of it. We have also seen many examples in the past, and a key purpose of securities and macroprudential regulations is to prevent such behaviour.
And finally, we have those aiming to manipulate the system for illegal purposes, like rogue traders, criminals, terrorists and nation-states. The inherent vulnerability of the financial system to adversarial attacks and the ease of manipulating AI engines creates potential for particular attack vectors that can be impossible to identify ex-ante. In other words, AI engine corruption becomes another form of cyber attack. With such an attack vector in place, an AI engine becomes a tool for well-resourced hostile agents that might be difficult or impossible to prevent in real time. That can facilitate nation-state attacks on critical infrastructure, especially since a superior ability to identify optimal timing strategies allows them to solve the problem of double coincidence, as discussed in Daníelsson et al. (2016), creating heightened system fragility, perhaps by manufacturing a liquidity crisis as an attack amplifier.
@@ 5.2 Misinformed use and overreliance on AI
The second channel emerges from humans using AI incorrectly. That is particularly relevant when AI is used in domains with unclear objectives, which is very common in the financial system. As events become less frequent and more serious, data-driven decisions increasingly involve interpretation, intuition and nuance, none of which AI in its current form is good at. As we discuss on Page 2.4, many crises have been resolved by very creative solutions that had never been used before.
@@
Because successful AI architectures are data-driven engines, their analysis is predominantly founded on statistical regularities. Such AI engines are designed to provide advice and responses to prompts, even if they have very low confidence about the accuracy of the answer. They can even make up facts or present arguments that sound plausible but would be considered flawed by an expert, both instances of the broader phenomenon of AI hallucination. The risk is that the AI engines will present confident recommendations about outcomes they know little about. To overcome that, the engines will have to provide assessment of the statistical accuracy of its recommendations. Here, it will be helpful if the authorities overcome their frequent reluctance to adopt consistent quantitative frameworks for measuring and reporting on the statistical accuracy of their data-based inputs and outputs.
Ultimately, these issues arise because of how AI creates trust. There are many examples where technology was initially met with scepticism. As it performs, technology is then increasingly accepted, as we have seen with AI landing passenger aeroplanes, AI-guided surgery and self-driving cars. Even if we only allow AI to execute the simplest tasks inside the financial authorities, as it starts performing at the same level or better than its human counterparts, but at a much lower cost, we can expect AI to become increasingly welcomed. In other words, its very success creates trust. As that trust is earned on relatively simple and safe repetitive tasks, and seeing AI succeed in more complex jobs, it will be asked to take on ever more sophisticated tasks.
As trust builds up, the critical risk is that we become so dependent on AI that the authorities cannot exercise control without it. In a sense, AI optimises to become irreplaceable. Therefore, the authorities can become dependent on an engine they do not fully understand or control and cannot turn off because that would have unpredictable consequences.
5.3 AI misalignment
@@
The third channel for how AI can be destabilising relates to difficulties in aligning the behaviour of AI engines with the objectives of their operators or those of the authorities regulating their private sector owners. The latter is particularly challenging if the objectives of financial regulations cannot be defined precisely, as usually is the case, as we discussed in section 2.4. Problems of misalignment arise when AI makes decisions because of the impossibility of fully specifying all objectives it has to meet, a form of incomplete contracting. The best solution an engine comes up with might be undesirable for the institution it works for. While such individual misbehaviour is certainly all too common in today’s human-centred financial institutions, it becomes more prevalent as AI use increases. Humans have been taught ethics and have general notions such as “don’t cheat, steal, no violence.” Most organisations try to recruit individuals for whom that education was successful. AI might require a similar broad education if its objectives cannot be fully specified ex-ante.
@@
Scheurer et al. (2023) provides an example of how individual AI can spontaneously choose to break regulations in its pursuit of profit. Using GPT-4 to analyse stock trading, they told their AI engine that insider trading was unacceptable. When they then gave the engine an illegal stock tip, it proceeded to trade on it and lie to the human overseers. Here, AI is simply engaging in the same type of illegal behaviour so many humans have done before.
@@
The problem of misalignment is particularly difficult when several AI engines interact, which can amplify existing fragilities. For example, market manipulation is most successful when several investors collude. Similarly, the payoff of many trading activities, such as carry trades and short sales, increases with the number of investors participating in them. A particularly damaging consequence of the strategic complementarities prevalent in the financial system relates to the objectives of financial institutions. They maximise profit most of the time, perhaps 999 days out of a thousand, but optimise for survival during stress. The behaviour on that one day out of a thousand is particularly damaging when financial institutions seek safety by withdrawing liquidity, which can lead to destabilising dynamics such as fire sales, bank runs, and credit crunches. The very high level of AI performance can, perhaps paradoxically, increase the likelihood of coordinated investor behaviour as private sector AI engines will excel at. They might find better strategies for exploiting strategic complementarities for private gain, at the risk of lowering market quality or even threatening financial stability.
@@
The most difficult problems of misalignment arise due to the interaction of the various AI working throughout the industry. Long before computers got involved with the financial system, humans coordinated damaging behaviour due to the misalignment of individual short-term incentives with the long-term incentives of the same market participants and society at large. The crises of 1763, 1914 and 1929 (see, e.g. Daníelsson, 2022) are just two of many such examples, as is the tulip mania in the Netherlands in the 1630s, see Aliber and Kindleberger (2015). The use of computers amplifies the potential for such undesirable outcomes as two recent examples illustrate. The largest stock market crash in history happened in 1987 when global markets fell by 23% in one day. The reason was trading algorithms that, in a way not foreseen, coordinated in creating a vicious feedback between prices collapsing and algorithm-induced selling, as documented by Gennotte and Leland (1990). Similarly, the quant crisis in the summer of 2007, as noted by Khandani and Lo (2007), was due to trading algorithms trained to sell on falling prices but not being told of the potential for aggregate feedback caused by trading algorithms in a number of funds coordinating on selling. Such factors have further been at the root of the various flash crashes. AI will likely amplify such problems of misalignment.
@@ 5.4 Resource concentration
The oligopolistic nature of the AI analytics business further strengthens that efficiency-induced procyclicality. ML design, input data and compute affect AI engines’ ability. These are controlled mainly by few technology and information companies, which continue to merge, creating an increasingly oligopolistic market. The median salary for specialists in data, analytics and artificial intelligence in US banks was \$901,000 in 2022 and \$676,000 in Europe. [^8] https://www.bloomberg.com/news/articles/2023-11-28/goldman-raided-by-recruiters-in-wall-street-fight-for-ai-talent Since there is a considerable shortage of the necessary human capital and the productivity of those experts is directly affected by the network effects of working with other experts, as well as the data and compute available to them, this alone puts designing the most effective AI engines out of the reach of all but the largest financial institutions. Furthermore, the currently largest neural networks have trillions of parameters, GPT 4 with 1.7 trillion, requiring significant funding and access to specialised data centres with the requisite GPUs. The current top engines have hundreds of millions of dollars in annual budgets. OpenAI reported that the cost of training GPT-4 exceeded \$100 million. [^9] https://www.wired.com/story/openai-ceo-sam-altman-the-age-of-giant-ai-models-is-already-over
Finally, financial data vendors have concentrated considerably over the past few years, with only a handful left, such as S&P Global, Bloomberg and LSEG. They have unified databases with standardised labelling and graph structures linking disparate categories of data, including proprietary and not publicly available data, facilitating training. It is a concern that we have not sufficiently appreciated the potential for AI-technology-data oligopolies in the recent wave of data vendor mergers.
Taken together, AI-driven financial analytics is characterised by increasing returns technologies with high entry costs, making an oligopolistic market structure highly likely. In such an environment, outsourcing analytics to one of those vendors might be the best, maybe even the only feasible solution for all but the largest financial companies. Anecdotal private sector evidence indicates that many financial institutions already prefer to outsource data-driven applications to specialists. A particular example is risk management as a service, RMaaS, with a leading vendor being BlackRock’s AI-fuelled Aladdin. The risk is harmonised beliefs and actions, driving procyclical behaviour and exposing users to the same blind spots.
5.5 AI evading control
That means AI, either autonomously or at human instruction, would likely be particularly good at finding strategies for avoiding both internal controls in a financial institution and regulatory oversight, making the two-sided principal-agent problem we discuss in Section 2.5 particularly challenging. In other words, the more sophisticated AI becomes, the more difficult it will be to exercise authority over it. After all, the controllers need to understand the AI objectives to design the correct incentives. And given the current state of interpretability of foundation models, this is very challenging.

### 03 Feb 2024 v4

@@ 5 The channels for how AI can destabilise the financial system
While we anticipate that AI will be broadly beneficial to the financial system, it may also increase systemic financial risk, defined as a serious financial crisis that has a high chance of causing an economic recession. The costs of systemic crises are very high, in the trillions of dollars for large economies, as noted by Barnichon et al. (2022). As such crises are usually caused by excessive low-quality lending, amplified by high leverage and asymmetric information, they should be easily preventable as macroprudential regulations aim to do. It is not so in reality. The almost infinite complexity of the financial system makes it very hard to design effective regulations that simultaneously prevent excesses and do not hamper economic growth too much. On top of that, because the economy is usually growing rapidly before a crisis, there is strong political opposition to any measures that might threaten that growth.
@@
We build on extant work on AI safety, like Weidinger et al. (2022), Bengio et al. (2023) and Shevlane et al. (2023), which identifies several societal risks arising from AI use, including malicious use, misinformation and loss of human control. When augmenting that, we propose four channels for how AI use may destabilise the financial system, all of which follow from the AI issues discussed in Section 3 and the conceptual challenges in Section 2.
@@ 5.1 Malicious use of AI
The first channel arises from how the operators of financial AI use them to optimise against the system for malicious purposes. Financial institutions and their staff deploy considerable resources in their pursuit of profit and usually are not very concerned about the wider social impact of their activities. They can change the system in ways that benefit them while not being detectable by others, as we have seen many times in the past. AI provides such actors with new opportunities, either via adversarial attacks by feeding particular data into its training algorithms or influencing network design. That might not even be necessary since it is straightforward to exploit for private, even illicit, gain gaps in how AI sees its responsibilities.
One way such malicious use might occur is by manipulating governance processes. The financial authorities and internal control aim to align the interests of system participants to the organisation that employs them and society. As AI is particularly effective in finding profitable loopholes and amplifying vulnerabilities, it can facilitate misbehaviour that, while legal, is damaging both to society and even the institution employing the AI.
The operator of the AI could take that one step further and use it to deliberately create market stress, which, of course, is profitable to those forewarned. We have seen many examples of such conduct in the past, and a key purpose of securities and macroprudential regulations is to prevent such behaviour.
And finally, we have those aiming to manipulate the system for illegal purposes, like rogue traders, criminals, terrorists and nation-states. The inherent vulnerability of the financial system to adversarial attacks and the ease of manipulating AI engines create potential for particular attack vectors that can be impossible to identify ex-ante. One example comes from Hubinger et al. (2024) who identify what they call “sleeper agent” that usually acts as expected but gives deceptive answers when given special instructions. Here, AI engine corruption becomes another form of cyber attack. With such an attack vector in place, an AI engine becomes a tool for well-resourced hostile agents that might be difficult or impossible to prevent in real time, facilitating nation-state attacks on critical infrastructure. That is particularly relevant when a superior ability to identify optimal timing strategies allows them to solve the problem of double coincidence, as discussed in Daníelsson et al. (2016), creating heightened system fragility, perhaps by manufacturing a liquidity crisis as an attack amplifier.
@@ 5.2 Misinformed use and overreliance on AI
The second channel emerges from humans using AI incorrectly. That is particularly relevant when AI is used in domains with unclear objectives, which is very common in the financial system. As events become less frequent and more serious, data-driven decisions increasingly involve extrapolation, interpretation, intuition and nuance, none of which AI in its current form is good at. As we discuss on Page 2.4, many crises have been resolved by very creative solutions that had never been used before.
@@
Because successful AI architectures are data-driven engines, their analysis is predominantly founded on statistical regularities. Such AI engines are designed to provide advice and responses to prompts, even if they have very low confidence about the accuracy of the answer. They can even make up facts or present arguments that sound plausible but would be considered flawed by an expert, both instances of the broader phenomenon of AI hallucination. The risk is that the AI engines will present confident recommendations about outcomes they know little about. To overcome that, the engines should be required to provide an assessment of the statistical accuracy of their recommendations. Here, it will be helpful if the authorities overcome their frequent reluctance to adopt consistent quantitative frameworks for measuring and reporting on the statistical accuracy of their data-based inputs and outputs.
5.3 AI misalignment and control avoidance
@@
The third channel for how AI can be destabilising relates to difficulties in aligning the behaviour of AI engines with the objectives of both their owners and financial authorities. The latter is particularly challenging if the objectives of financial regulations cannot be defined precisely, as usually is the case, as we discussed in Section 2.4. Problems of misalignment arise when AI makes decisions because of the impossibility of fully specifying all objectives it has to meet, a form of incomplete contracting. The best solution an engine comes up with might be undesirable for the institution it works for. While such individual misbehaviour is certainly all too common in today’s human-centred financial institutions, it becomes more prevalent as AI use increases. Humans have been taught ethics and have general notions such as “don’t cheat, steal, no violence.” Most organisations try to recruit individuals for whom that education was successful. AI might require a similar broad education if its objectives cannot be fully specified ex-ante.
@@
Scheurer et al. (2023) provides an example of how individual AI can spontaneously choose to break regulations in their pursuit of profit. Using GPT-4 to analyse stock trading, they told their AI engine that insider trading was unacceptable. When they then gave the engine an illegal stock tip, it proceeded to trade on it and lie to the human overseers. Here, AI is simply engaging in the same type of illegal behaviour so many humans have done before.
@@
The problem of misalignment is particularly difficult when several AI engines interact. For example, market manipulation is most successful when several investors collude. Similarly, the payoff of many trading activities, such as carry trades and short sales, increases with the number of investors participating in them. A particularly damaging consequence of the strategic complementarities prevalent in the financial system relates to the objectives of financial institutions. They maximise profit most of the time, perhaps 999 days out of a thousand, but optimise for survival during stress. The behaviour on that one day out of a thousand is particularly damaging when financial institutions seek safety by withdrawing liquidity, which can lead to destabilising dynamics such as fire sales, bank runs, and credit crunches. The very high level of AI performance can, perhaps paradoxically, increase the likelihood of damaging coordinated behaviour. They might find better strategies for exploiting strategic complementarities for private gain, at the risk of lowering market quality or even threatening financial stability.
@@
The most difficult problems of misalignment arise due to the interaction of the various AI working throughout the industry. Long before computers got involved with the financial system, humans coordinated damaging behaviour due to the misalignment of individual short-term incentives with the long-term incentives of the same market participants and society at large. The crises of 1763, 1914 and 1929 (see, e.g. Daníelsson, 2022) are just two of many such examples, as is the tulip mania in the Netherlands in the 1630s, see Aliber and Kindleberger (2015). The use of computers amplifies the potential for such undesirable outcomes as two recent examples illustrate. The largest stock market crash in history happened in 1987 when global markets fell by 23% in one day. The reason was trading algorithms that, in a way not foreseen, coordinated in creating vicious feedback between prices collapsing and algorithm-induced selling, as documented by Gennotte and Leland (1990). Similarly, the quant crisis in the summer of 2007, as noted by Khandani and Lo (2007), was due to trading algorithms trained to sell on falling prices but not being told of the potential for aggregate feedback caused by trading algorithms in a number of funds coordinating on selling. Such factors have further been at the root of the various flash crashes. AI will likely amplify such problems of misalignment.
That means AI, either autonomously or at human instruction, would likely be particularly good at finding strategies for avoiding both internal controls in a financial institution and regulatory oversight, making the two-sided principal-agent problem we discuss in Section 2.5 particularly challenging. In other words, the more sophisticated AI become, the more difficult it will be to exercise authority over them. After all, the controllers need to understand the AI objectives to design the correct incentives. And given the current state of interpretability of foundation models, this is very challenging.
@@ 5.4 Resource concentration
The oligopolistic nature of the AI analytics business further strengthens that efficiency-induced procyclicality. ML design, input data and compute affect AI engines’ ability. These are controlled mainly by a few technology and information companies, which continue to merge, creating an increasingly oligopolistic market. The median salary for specialists in data, analytics and artificial intelligence in US banks was \$901,000 in 2022 and \$676,000 in Europe. [^8] https://www.bloomberg.com/news/articles/2023-11-28/goldman-raided-by-recruiters-in-wall-street-fight-for-ai-talent Since there is a considerable shortage of the necessary human capital and the productivity of those experts is directly affected by the network effects of working with other experts, as well as the data and compute available to them, this alone puts designing the most effective AI engines out of the reach of all but the largest financial institutions. Furthermore, the currently largest neural networks have trillions of parameters, GPT 4 with 1.7 trillion, requiring significant funding and access to specialised data centres with the requisite GPUs. The current top engines have hundreds of millions of dollars in annual budgets. OpenAI reported that the cost of training GPT-4 exceeded \$100 million. [^9] https://www.wired.com/story/openai-ceo-sam-altman-the-age-of-giant-ai-models-is-already-over
Finally, financial data vendors have concentrated considerably over the past few years, with only a handful left, such as S&P Global, Bloomberg and LSEG. They have unified databases with standardised labelling and graph structures linking disparate categories of data, including proprietary and not publicly available data, facilitating training. It is a concern that neither the competition nor the financial authorities appear to have fully appreciated the potential for increased systemic risk due to oligopolistic AI technology in the recent wave of data vendor mergers.
Taken together, AI-driven financial analytics is characterised by increasing returns to scale technologies with high entry costs, making an oligopolistic market structure highly likely. In such an environment, outsourcing analytics to one of those vendors might be the best, maybe even the only feasible solution for all but the largest financial companies. Anecdotal private sector evidence indicates that many financial institutions already prefer to outsource data-driven applications to specialists. A particular example is risk management as a service, RMaaS, with a leading vendor being BlackRock’s AI-fuelled Aladdin. The risk is harmonised beliefs and actions, driving procyclical behaviour and exposing users to the same blind spots.